OpenStudy (anonymous):

If you can answer this, you are a genius! 3.Ellen works for a high-speed rail company that wants to develop a new rail line. Ellen’s project is to find a train that is the second fastest in the world. The Shinkansen Bullet Train in Japan is reported to go as fast as 320 kilometers per hour. The TGV train in France can reach speeds of 89.44 meters per second. Explain to Ellen how to find what speed her new train must go to be the second fastest in this group. Then, find an appropriate speed for Ellen’s train in miles per hour and use the formula distance = speed • time to find how far the tra

OpenStudy (anonymous):

in can go in 2.5 hours. good luck..

OpenStudy (ranga):

This question is all about unit conversion. There are 1000 meters in a Kilometer There are 3600 seconds in one hour 1mi = 1.6093km Convert everything to mph. Can you take it from here?

OpenStudy (anonymous):

Not at all, could you explain a bit more?

OpenStudy (ranga):

Convert all speed to mi / hr Japanese train is 320 km / hr 1mi = 1.6093km or 1 km = 1/1.6093 mi So 320 km / hr = (320) (1/1.6093) mi / hr = 198.8 mi / hr French train is 89.44 meters per second. Convert it first to meter per hour by multiplying by 3600 as there are 3600 seconds in an hour 89.44 x 3600 = 321984 meters per hour Convert meters to km by dividing by 1000 as there are 1000 meters in a kilometer 321984 / 1000 = 321.984 km / hr convert to miles per hour by dividing by 1.6093 321.984 / 1.6093 = 200.1 mi / hr So the fastest train is the French train at 200.1 mi / hr The Japanese train is at 198.8 mi / hr To be the second fastest train, the speed has to be between those two numbers

OpenStudy (ranga):

I will take a midpoint (or average) between those two numbers: Ellen's train must be at speed (200.1 + 198.8)/2 = 199.5 mi / hr to be the second fastest train. In 2.5 hours the train will travel: 199.5 x 2.5 = 499 miles.

OpenStudy (anonymous):

Wow, that actually makes sense now. It also doesn't look like a foreign language! haha thank you!!

OpenStudy (ranga):

you are welcome.

OpenStudy (anonymous):

Maybe you should do the work yourself? Ranga it doesn't seem you are helping...just helping her to cheat

OpenStudy (anonymous):

it sure as heck helped me. Its not cheating when you truly dont understand how to set up a problem. I appreciate the help Ranga!!

OpenStudy (anonymous):

shortyshay..you know its cheating... I'm sorry you don't understand school is to learn and not get someone else to do your work for you